Riverbed Limited reported profit of $412,580 for its November 30, 2021, year end. Cash dividends of $74,600 on the common shares and $69,200 on the noncumulative preferred shares were declared and paid during the year. The following information is availab

Dec. 1, 2020
The opening number of common shares was 61,000.
Feb. 28, 2021
Sold 9,200 common shares for $202,400 cash.
May 31, 2021
Reacquired 4,600 common shares for $92,000 cash.
Nov. 1, 2021
Issued 13,800 common shares in exchange for land with a fair value of $300,000.



Calculate the weighted average number of common shares for the year.
Answer #1

To determine the weighted average number of common shares outstanding, we need to consider the shares outstanding for each period based on the given information:

  1. Opening shares (December 1, 2020): 61,000 shares

  2. Shares sold on February 28, 2021: 9,200 shares

  3. Shares reacquired on May 31, 2021: 4,600 shares

  4. Shares issued on November 1, 2021: 13,800 shares

To calculate the weighted average number of common shares outstanding, we need to assign weights to each period based on the portion of the year they were outstanding.

Weighted average number of shares = (Number of shares × Number of days outstanding) / Total number of days in the year

Here's the calculation for each period:

  1. Opening shares (December 1, 2020 - February 28, 2021): Number of shares = 61,000 Number of days outstanding = 90 days (December 1, 2020 - February 28, 2021) Total number of days in the year = 365 days Weighted average shares for this period = (61,000 × 90) / 365 = 15,082 shares

  2. Shares sold on February 28, 2021: Number of shares = 9,200 Number of days outstanding = 274 days (February 28, 2021 - November 30, 2021) Total number of days in the year = 365 days Weighted average shares for this period = (9,200 × 274) / 365 = 6,877 shares

  3. Shares reacquired on May 31, 2021: Number of shares = -4,600 (negative sign indicates shares reacquired) Number of days outstanding = 213 days (May 31, 2021 - November 30, 2021) Total number of days in the year = 365 days Weighted average shares for this period = (-4,600 × 213) / 365 = -2,685 shares

  4. Shares issued on November 1, 2021: Number of shares = 13,800 Number of days outstanding = 30 days (November 1, 2021 - November 30, 2021) Total number of days in the year = 365 days Weighted average shares for this period = (13,800 × 30) / 365 = 1,138 shares

Now, we can calculate the overall weighted average number of common shares outstanding:

Weighted average number of common shares outstanding = Sum of all weighted average shares = 15,082 + 6,877 - 2,685 + 1,138 = 20,412 shares

Therefore, the weighted average number of common shares outstanding for Riverbed Limited is 20,412 shares.


answered by: mervetokaz